# Automata Learning from Preference and Equivalence Queries

### v2

@@ Introduction
Average Policy Regret  We employ Q-learning with counterfactual experiences for reward machines (CRM) (Toro Icarte et al. 2022) to obtain optimal policies for ground truth and learned RMs. We measured the empirical expected return of optimal policies learned from each type of RM. Average regret for a given task was measured as the difference between the empirical return under the ground truth RM for that task (averaged over 100 CRM trials) and the empirical return under the learned RM (with 10 CRM trials per learned RM, then averaging all $100\times 10=1000$ trials). Regret goes to 0 as the number of samples tested by the teacher per equivalence query increases (Figure 3, right column). The Appendix A describes regret computation details.
@@
Figure 4 shows query complexity results. We measure preference query complexity of exact and PAC-ID Remap, as a function of the number of unique sequences stored in the table upon termination. Exact Remap (upper left) displays a trendline of $C=0.2121N\ln N$ with $R^{2}=0.9963$, where $C$ is the number of queries, and $N$ is the number of unique sequences in the observation table. PAC-ID Remap (upper right) tends to make significantly more preference queries about unique sequences compared to exact Remap due to the sampling process. However, the number of preference queries is still $\mathcal{O}(N\ln N)$ due to randomized quicksort and linear merge comparison complexity.
@@
The termination phase diagram in Figure 4 illustrates the taxi distance from $(0,1)$ to $(|\Sigma^{O}|,n)$ is the maximum number of equivalence queries Remap makes (Theorem 3). Progress (Lemma 1 and Theorem 2) towards termination (Corollary 1) occurs whenever a new hypothesis is made. Remap can terminate early when all variables have correct values and the required number of states is reached.
Related Work
@@
Active approaches for learning automata are variations or improvements of Angluin’s seminal $\textrm{L}^{*}$ algorithm (1987), featuring membership and equivalence queries. We consider an alternative formulation: actively learning automata from preference and equivalence queries featuring feedback. We first discuss adaptations of $\textrm{L}^{*}$ for learning variants of finite automata, including reward machine specific variants.
@@
Learning Finite Automata  Angluin (1987) introduced $\textrm{L}^{*}$ to learn deterministic finite automata (DFAs). Remap has similar theoretical guarantees as $\textrm{L}^{*}$, but utilizes a symbolic observation table, rather than an evidence-based one. Other algorithms adopt the evidence-based table of $\textrm{L}^{*}$ to learn: symbolic automata (Drews and D’antoni 2017; Argyros and D’antoni 2018), where Boolean predicates summarize state transitions; weighted automata (Bergadano and Varricchio 1994; Balle and Mohri 2015) which feature valuation semantics for sequences on non-deterministic automata; probabilistic DFAs (Weiss, Goldberg, and Yahav 2019), a weighted automata that models distributions of sequences. None of these approaches uses preference queries.
@@
However, Shah et al. (2023) considers active, cost-based selection between membership and preference queries to learn DFAs, relying on a satisfiability encoding of the problem. They assume a fixed hypothesis space and have probabilistic guarantees for termination and correctness. Remap, through unification, navigates a sequence of hypothesis spaces, each guaranteed to contain a concrete hypothesis satisfying current constraints, and has theoretical guarantees of correctness, minimalism, and termination under exact and PAC–identification settings.
@@
Furthermore, learning finite automata from preference information relates to the novel problem of learning reward machines (Icarte et al. 2018) from preferences. Learning Markovian reward functions from preferences has be studied extensively using neural (Bıyık, Talati, and Sadigh 2022; Sadigh et al. 2017) and interpretable decision tree (Bewley and Lécué 2021; Kalra and Brown 2023) representations, but approaches for learning reward machines primarily adapt evidence-based finite automata learning approaches.
@@
Reward Machine Variants  Several reward machine (RM) variants have been proposed. Classical RMs (Icarte et al. 2018; Toro Icarte et al. 2019) have deterministic transitions and rewards; probabilistic RMs (Dohmen et al. 2022) model probabilistic transitions and deterministic rewards; and stochastic RMs (Corazza, Gavran, and Neider 2022) pair deterministic transitions with stochastic rewards. Symbolic RMs (Zhou and Li 2022) are deterministic like classical RMs, but feature symbolic reward values in place of concrete values. Zhou and Li (2022) apply Bayesian inverse reinforcement learning (BIRL) to infer optimal reward values and actualize symbolic RMs into classical RMs, and require a symbolic RM sketch. Remap requires no sketch, since it navigates over a hypothesis space of symbolic RMs and outputs a concrete classical RM upon termination.
@@
Learning Reward Machines  Many RM learning algorithms assume access to explicit reward samples via environment interaction. Given a maximum RM size, Toro Icarte et al. (2019) apply discrete optimization to arrive at a perfect classical RM. Xu et al. (2020) learn a minimal classical RM by combining regular positive negative inference (Dupont 1994) with Q-learning for RMs (Icarte et al. 2018), and apply constraint solving to ensure each hypothesis RM is consistent with observed reward samples. Corazza et al. (2022) extended the method to learn stochastic RMs. Topper et al. (2024) extends BIRL to learn classical RMs using simulated annealing, but needs the number of states to be supplied, and requires empirical tuning of hyperparameters. $\textrm{L}^{*}$ based approaches have also been used to learn classical (Tappler et al. 2019; Gaon and Brafman 2020; Xu et al. 2021) and probabilistic (Dohmen et al. 2022) RMs, relying on concrete observation tables. Gaon and Brafman and Xu et al. use a binary observation table, while Tappler et al. and Dohmen et al. record empirical reward distribution table entries.
@@
Conclusion
We introduce the problem of learning Moore machines from preferences and propose Remap, an $\textrm{L}^{*}$ based algorithm, wherein a strong learner with access to a constraint solver is paired with a weak teacher capable of answering preference queries and providing counterexample feedback in the form of a constraint. Unification applied to a symbolic observation table permits symbolic hypothesis space navigation; the constraint solver enables concrete hypotheses. Remap has theoretical guarantees for correctness, termination, and minimalism under both exact and PAC–identification settings, and it has been empirically verified under both settings when applied to learning reward machines. Next steps include investigating more realistic preference models, handling variable strength feedback, and handling inconsistency.
Appendix A Technical Appendix
@@
Experimental Details
Hardware
Experiments were run on a server with 512GB of memory and 4 Intel(R) Xeon(R) Gold 6248R CPUs. Each experimental run was executed on a single core. The operating system was Ubuntu 22.04.4. The relevant software libraries are listed in the code appendix remap/README file.
OfficeWorld and CraftWorld Domains
@@
Handling Reward Machine Incompleteness.
The classical reward machines specified by Icarte et al. (2018) in the OfficeWorld and CraftWorld domains are incomplete automata, in that not all states have transitions defined. Specifically, classical reward machines have a terminal state from which no transitions can occur. However, an incomplete reward machine can be converted into a complete one by adding transitions from all terminal states to a single, special absorbing “HALT” state.
@@
Handling Reward Machine Nondeterminism.
Classical reward machines have a deterministic definition. However, some of the classical reward machines specified by Icarte et al. (2018) are nondeterministic, in that from a given state, multiple transitions can be satisfied using under a given set of true propositions. This was the case for the reward machines for tasks 5 through 10 in the CraftWorld domain.
@@
Test Set Generation for Classification Accuracy
@@
Regret Computation Details
@@
We utilized Q-learning with counterfactual experiences for reward machines (CRM) (Toro Icarte et al. 2022) to obtain optimal policies for each reward machine. We set the discount factor to $\gamma=0.9$, and for the OfficeWorld reward machines, each policy was trained for a total of $2\times 10^{5}$ steps, and $2\times 10^{6}$ steps for CraftWorld reward machines. Each reward machine was trained for at least 10 seeds. Observing the return curves, we concluded that by $1\times 10^{5}$ steps, the policy was already optimal for OfficeWorld domains, and by $1\times 10^{6}$ steps for CraftWorld domains. We compute the average reward per step of the optimal policy by summing the total return of the policy over the last $10^{5}$ (OfficeWorld) or $10^{6}$ (CraftWorld) steps:
@@
Example Comparing REMAP and $\textrm{L}^{*}$
Figure 2 of the main paper contained an abridged example comparing REMAP and $\textrm{L}^{*}$. We show the full example in Figure 6 of the Appendix showing how REMAP and $\textrm{L}^{*}$ learn the Moore machine shown in part (l).
@@
Additional Scaling Measurements
@@
We present a full set of termination phase diagrams for learning the CraftWorld reward machines for tasks 5 through 10 in Figure 8. Columns correspond to tasks, while rows correspond to number of samples the teacher make per equivalence query. Each plot contains 100 paths through the termination phase space, where the $x$-axis is the number of known representative values, and the $y$-axis is the number of states in the hypothesis. Each node along the path corresponds with an event in Remap. The start of each path always starts at $(0,1)$, since there is always an initial state, but no known representatives. Green $X$’s represent when an equivalence query is made. Blue circles represent the immediate result of a closure operation, while orange squares represent the immediate result of a consistency operation. The red star represents the upper bound on number of states and number of known representative values. Observe that between consecutive equivalence queries, at least one of the number of states or the number of known representative values must increase: in particular, if the number of known representative values does not increase as the result of an equivalence query, the number of states must increase. Additionally, it is possible for Remap to terminate early, when the number of states has reached the upper bound, and when the satisfying solution to the constraints has all correct values.
@@
Figure 8: All phase diagram termination plots for CraftWorld tasks 5 through 10. From left to right: Task 5, 6, 7, 8, 9, 10. From top to bottom: Teacher tests 25, 50, 100, 200, 300, 400, 500 samples per equivalence query. The X-axis is number of known representative values in the hypothesis, the Y-axis is the number of states in the hypothesis. Legend: blue circle is a closure operation, orange square is a consistency operation, green x is an equivalence query, and the red star represents the the upper bound termination condition. We observe it is possible to terminate early, prior to reaching the upper bound. Each plot contains 100 individual paths through phase space.
Remap Algorithms
@@
Technical Proofs
We first establish the correctness of constructing a hypothesis deterministic finite automaton from a symbolic observation table via Algorithm 7, in Theorem 1 via lemmas 1-5. The proof strategies for lemmas 3 through 5 generally follows that of Angluin (1987), with appropriate adjustments to account for a set of possible hypotheses, compared to just a single hypothesis.
Lemma 1.
@@
MakeHypothesis$(\langle S,E,T;\mathcal{C},\Gamma\rangle)$. All hypotheses in $\mathcal{H}$ are structurally identical. Two hypotheses $h_{1}$ and $h_{2}$ are structurally identical if they have equivalent states and equivalent transitions.
Proof.
We first show that all hypotheses in $\mathcal{H}$ must be structurally identical by construction, but have different $\hat{L}$ functions. Identical structure for all hypotheses in $\mathcal{H}$ can be shown by observing that lines 1-3 of Algorithm 7 are the same for each possible hypothesis in $\mathcal{H}$ for a given $\langle S,E,T;\mathcal{C},\Gamma\rangle$ input. Therefore, the only differences in the output hypothesis are due to lines 4 and 5. Since lines 1-3 construct the set of states and the transition function for each hypothesis, each hypothesis in $\mathcal{H}$ must therefore be structurally identical to all other hypotheses in $\mathcal{H}$. Lines 4 and 5 together select a satisfying solution for the free variables in $\langle S,E,T;\mathcal{C},\Gamma\rangle$, subject to the constraints $\mathcal{C}$. Therefore, each unique hypothesis in $\mathcal{H}$ is corresponds to a unique satisfying solution; the mapping between unique satisfying solutions and unique hypotheses is bijective. ∎
Lemma 2.
States in a Moore machine are represented by sets of sequences.
Proof.
Suppose we have a finite alphabet $\Sigma$, and a Moore machine, with set of states $Q$, initial state $q_{0}\in Q$, and transition function $\delta:Q\times\Sigma\rightarrow Q$. We define the initial state $q_{0}$ of the Moore machine to correspond with the empty sequence $\varepsilon$ which has length 00, so in fact we have $q_{0}=\delta(q_{0},\varepsilon)$, and for any $q\in Q$, $q=\delta(q,\varepsilon)$. The transition function can be extended recursively to $\delta:Q\times\left(\Sigma^{*}\right)\rightarrow Q$ by observing that $\delta(q_{0},\sigma\cdot\omega)=\delta(\delta(q_{0},\sigma),\omega)$, where $\sigma$ is a sequence of length of 1, and $\omega$ is a sequence of length of at least $1$. Similarly, $\delta(q_{0},\omega\cdot\sigma)=\delta(\delta(q_{0},\omega),\sigma)$. Based on this extended transition function, we can now consider the set $Q^{\prime}=\{\delta(q_{0},\sigma)|\sigma\in\Sigma\}$ to be a subset of $Q$.
@@
Proof.
We utilize the inductive proof of Angluin (1987). For the case of $s=\varepsilon$, with length 00, we have $\delta(q_{0},\varepsilon)=q_{0}=\textbf{row}(\varepsilon)$, which is true by definition. Now, let us assume that $\delta(q_{0},s)=\textbf{row}(s)$ holds for all $s\in(S\cup(S\cdot\Sigma))$ with lengths no greater than $k$. Let $s^{\prime}\in(S\cup(S\cdot\Sigma))$ be a sequence of length $k+1$ such that $=s\cdot\sigma$. If $s^{\prime}\in S$, then $s\in S$ because $S$ is prefix closed. If $s^{\prime}\in S\cdot\Sigma$, then $s\in S$. We can then show that $\delta(q_{0},s^{\prime})=\delta(q_{0},s\cdot\sigma)=\delta(\delta(q_{0},s),\sigma)=\delta(\textbf{row}(s),\sigma)=\textbf{row}(s\cdot\sigma)=\textbf{row}(s^{\prime})$. ∎
Lemma 4.
The entries of a unified, closed, and consistent symbolic observation table correspond to sequence classification consistent with constraints $\mathcal{C}$. Specifically, for a unified, closed, and consistent $\langle S,E,T;\mathcal{C},\Gamma\rangle$, let $\Sigma^{*}$ be partitioned into at most $k\leq|\Sigma^{O}|$ disjoint sets $F_{1},F_{2},...,F_{k}$, and let $\Sigma^{O}_{k}$ be a specific subset of $k$ distinct elements of $\Sigma^{O}$. A sequence $s\in\Sigma^{*}$ is a member of $F_{j}$ if and only if it is classified as $\sigma_{j}\in\Sigma^{O}_{k}$. Then hypothesis MakeHypothesis$(\langle S,E,T;\mathcal{C},\Gamma\rangle)$ satisfies $\delta(q_{0},s\cdot e)\in F_{j}$ if and only if $T(s\cdot e)=\sigma_{j}$ for every $s\in(S\cup(S\cdot\Sigma))\text{ and every }e\in E$.
Proof.
We adapt the inductive proof of Angluin (1987), but make appropriate adjustments for sequence classification. Let $s\in(S\cup(S\cdot\Sigma))$ and $e=\varepsilon$ be the base case. Clearly, $\delta(q_{0},s\cdot\varepsilon)=\delta(q_{0},s)=\textbf{row}(s)$ as shown previously. If $s\in S$, then $\textbf{row}(s)\in F_{j}$ if and only if $T(s)=\sigma_{j}$. If $s\in S\cdot\Sigma$, then $\textbf{row}(s)\in\hat{Q}$, since the symbolic observation table is closed. A $\textbf{row}(s^{\prime})\in\hat{Q}$ is in $F_{j}$ if and only if $T(s^{\prime})=\sigma_{j}$.

### v3

@@ 7.1 PAC–Identification Correctness Experiments
Average Policy Regret  We employ Q-learning with counterfactual experiences for reward machines (CRM) [41] to obtain optimal policies for ground truth and learned RMs. We measured the empirical expected return of optimal policies learned from each type of RM. Average regret for a given task was measured as the difference between the empirical return under the ground truth RM for that task (averaged over 100 CRM trials) and the empirical return under the learned RM (with 10 CRM trials per learned RM, then averaging all $100\times 10=1000$ trials). Regret goes to 0 as the number of samples tested by the teacher per equivalence query increases (Figure 4, right column). The Appendix 0.A.1.6 describes regret computation details.
@@ 7.2 Scaling Experiments
Figure 5 shows query complexity results. We measure preference query complexity of exact and PAC-ID Remap, as a function of the number of unique sequences stored in the table upon termination. Exact Remap (upper left) displays a trendline of $C=0.2114N\ln N$ with $R^{2}=0.99268$, where $C$ is the number of queries, and $N$ is the number of unique sequences in the observation table. PAC-ID Remap (upper right) tends to make significantly more preference queries about unique sequences compared to exact Remap due to the sampling process. However, the number of preference queries is still $\mathcal{O}(N\ln N)$ due to randomized quicksort and linear merge comparison complexity. In comparison, an $\textrm{L}^{*}$-based approach would use exactly $N$ membership queries (exactly linear in $N$ with a coefficient of $1$).
@@
The maximum number of equivalence queries Remap makes (Theorem 3) is the taxi distance from $(0,1)$ to $(|\Sigma^{O}|,n)$ in the termination phase diagram of Figure 5. Progress (Lemma 1 and Theorem 2) towards termination (Corollary 1) occurs whenever a new hypothesis is made. Remap can terminate early when all variables have correct values and the required number of states is reached.
8 Related Work
@@
Active approaches for learning automata are variations or improvements of Angluin’s seminal $\textrm{L}^{*}$ algorithm [5], featuring membership and equivalence queries. We consider an alternative formulation: actively learning automata from preference and equivalence queries featuring feedback. We first discuss adaptations of $\textrm{L}^{*}$ for learning variants of finite automata, including reward machine variants.
@@
Learning Finite Automata  Angluin [5] introduced $\textrm{L}^{*}$ to learn deterministic finite automata (DFAs). Remap has similar theoretical guarantees as $\textrm{L}^{*}$, but utilizes a symbolic observation table, rather than an evidence-based one. Other algorithms adopt the evidence-based table of $\textrm{L}^{*}$ to learn: symbolic automata [17, 7], where Boolean predicates summarize state transitions; weighted automata [9, 8] which feature valuation semantics for sequences on non-deterministic automata; probabilistic DFAs [44], a weighted automata that models distributions of sequences. None of these approaches uses preference queries.
@@
However, Shah et al. [38] considers active, cost-based selection between membership and preference queries to learn DFAs, relying on a satisfiability encoding of the problem. They assume a fixed hypothesis space and have probabilistic guarantees for termination and correctness. Remap, through unification, navigates a sequence of hypothesis spaces, each guaranteed to contain a concrete hypothesis satisfying current constraints, and has theoretical guarantees of correctness, minimalism, and termination under exact and PAC–identification settings.
@@
Furthermore, learning finite automata from preference information relates to the novel problem of learning reward machines [25] from preferences. Learning Markovian reward functions from preferences has be studied extensively using neural [11, 36] and interpretable decision tree [10, 27] representations, but approaches for learning reward machines primarily adapt evidence-based finite automata learning approaches.
@@
Reward Machine Variants  Several reward machine (RM) variants have been proposed. Classical RMs [25, 42] have deterministic transitions and rewards; probabilistic RMs [16] model probabilistic transitions and deterministic rewards; and stochastic RMs [14] pair deterministic transitions with stochastic rewards. Symbolic RMs [47] are deterministic like classical RMs, but feature symbolic reward values in place of concrete values. Zhou and Li [47] apply Bayesian inverse reinforcement learning (BIRL) to infer optimal reward values and actualize symbolic RMs into classical RMs, and require a symbolic RM sketch. Remap requires no sketch, since it navigates over a hypothesis space of symbolic RMs and outputs a concrete classical RM upon termination.
@@
Learning Reward Machines  Many RM learning algorithms assume access to explicit reward samples via environment interaction. Given a maximum RM size, Toro Icarte et al. [42] apply discrete optimization to arrive at a perfect classical RM. Xu et al. [45] learn a minimal classical RM by combining regular positive negative inference [18] with Q-learning for RMs [25], and apply constraint solving to ensure each hypothesis RM is consistent with observed reward samples. Corazza et al. [14] extended the method to learn stochastic RMs. Topper et al. [40] extends BIRL to learn classical RMs using simulated annealing, but needs the number of states to be supplied, and requires empirical tuning of hyperparameters. $\textrm{L}^{*}$ based approaches have also been used to learn classical [39, 20, 46] and probabilistic [16] RMs, relying on concrete observation tables. Gaon and Brafman [20] and Xu et al. [46] use a binary observation table, while Tappler et al. [39] and Dohmen et al. [16] record empirical reward distribution table entries.
@@
9 Conclusion
We introduce the problem of learning Moore machines from preferences and propose Remap, an $\textrm{L}^{*}$ based algorithm, wherein a strong learner with access to a constraint solver is paired with a weak teacher capable of answering preference queries and providing counterexample feedback in the form of a constraint. Unification applied to a symbolic observation table permits symbolic hypothesis space navigation; the constraint solver enables concrete hypotheses. Remap has theoretical guarantees for correctness, termination, and minimalism under both exact and PAC–identification settings, and it has been empirically verified under both settings when applied to learning reward machines. Future work will expound on more realistic preference models, variable strength feedback, and inconsistency.
{credits}
[…]
The authors have no competing interests to declare that are relevant to the content of this article.
Appendix 0.A Technical Appendix
@@
0.A.1 Experimental Details
0.A.1.1 Hardware
Experiments were run on a server with 512GB of memory and 2 Intel(R) Xeon(R) Gold 6248R CPUs. Each experimental run was executed on a single core. The operating system was Ubuntu 22.04.4. The relevant software libraries are listed in the code appendix remap/README file.
0.A.1.2 OfficeWorld and CraftWorld Domains
@@
0.A.1.3 Handling Reward Machine Incompleteness.
The classical reward machines specified by Icarte et al. [25] in the OfficeWorld and CraftWorld domains are incomplete automata, in that not all states have transitions defined. Specifically, classical reward machines have a terminal state from which no transitions can occur. However, an incomplete reward machine can be converted into a complete one by adding transitions from all terminal states to a single, special absorbing “HALT” state.
@@
0.A.1.4 Handling Reward Machine Nondeterminism.
Classical reward machines have a deterministic definition. However, some of the classical reward machines specified by Icarte et al. [25] are nondeterministic, in that from a given state, multiple transitions can be satisfied using under a given set of true propositions. This was the case for the reward machines for tasks 5 through 10 in the CraftWorld domain.
@@
0.A.1.5 Test Set Generation for Classification Accuracy
@@
0.A.1.6 Regret Computation Details
@@
We utilized Q-learning with counterfactual experiences for reward machines (CRM) [41] to obtain optimal policies for each reward machine. We set the discount factor to $\gamma=0.9$, and for the OfficeWorld reward machines, each policy was trained for a total of $2\times 10^{5}$ steps, and $2\times 10^{6}$ steps for CraftWorld reward machines. Each reward machine was trained for at least 10 seeds. Observing the return curves, we concluded that by $1\times 10^{5}$ steps, the policy was already optimal for OfficeWorld domains, and by $1\times 10^{6}$ steps for CraftWorld domains. We compute the average reward per step of the optimal policy by summing the total return of the policy over the last $10^{5}$ (OfficeWorld) or $10^{6}$ (CraftWorld) steps:
@@
0.A.2 Example Comparing REMAP and $\textrm{L}^{*}$
Figures 2 and 3 of the main paper contained an abridged example comparing REMAP and $\textrm{L}^{*}$. We show the full example in Figure 7 of the Appendix showing how REMAP and $\textrm{L}^{*}$ learn the Moore machine shown in part (l).
@@ 0.A.2 Example Comparing REMAP and $\textrm{L}^{*}$
0.A.3 Additional Scaling Measurements
@@ 0.A.4 Termination Plots
We present a full set of termination phase diagrams for learning the CraftWorld reward machines for tasks 5 through 10 in Figures 9, 10, 11, and 12. Columns correspond to tasks, while rows correspond to number of samples the teacher make per equivalence query. Each plot contains 100 paths through the termination phase space, where the $x$-axis is the number of known representative values, and the $y$-axis is the number of states in the hypothesis. Each node along the path corresponds with an event in Remap. The start of each path always starts at $(0,1)$, since there is always an initial state, but no known representatives. Green $X$’s represent when an equivalence query is made. Blue circles represent the immediate result of a closure operation, while orange squares represent the immediate result of a consistency operation. The red star represents the upper bound on number of states and number of known representative values. Observe that between consecutive equivalence queries, at least one of the number of states or the number of known representative values must increase: in particular, if the number of known representative values does not increase as the result of an equivalence query, the number of states must increase. Additionally, it is possible for Remap to terminate early, when the number of states has reached the upper bound, and when the satisfying solution to the constraints has all correct values.
  Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
@@
Figure 9: Phase diagram termination plots for CraftWorld tasks 5 through 10. Teacher tests 25 (top section) and 50 (bottom section) samples per equivalence query. Within each section: left to right (top row) Task 5, 6, 7; left to right (bottom row) Task 8, 9, 10. The X-axis is number of known representative values in the hypothesis, the Y-axis is the number of states in the hypothesis. Legend: blue circle is a closure operation, orange square is a consistency operation, green x is an equivalence query, and the red star represents the the upper bound termination condition. We observe it is possible to terminate early, prior to reaching the upper bound. Each plot contains 100 individual paths through phase space.
  Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 10: Phase diagram termination plots for CraftWorld tasks 5 through 10. Teacher tests 100 (top section) and 200 (bottom section) samples per equivalence query. Within each section: left to right (top row) Task 5, 6, 7; left to right (bottom row) Task 8, 9, 10. The X-axis is number of known representative values in the hypothesis, the Y-axis is the number of states in the hypothesis. Legend: blue circle is a closure operation, orange square is a consistency operation, green x is an equivalence query, and the red star represents the the upper bound termination condition. We observe it is possible to terminate early, prior to reaching the upper bound. Each plot contains 100 individual paths through phase space.
  Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 11: Phase diagram termination plots for CraftWorld tasks 5 through 10. Teacher tests 300 (top section) and 400 (bottom section) samples per equivalence query. Within each section: left to right (top row) Task 5, 6, 7; left to right (bottom row) Task 8, 9, 10. The X-axis is number of known representative values in the hypothesis, the Y-axis is the number of states in the hypothesis. Legend: blue circle is a closure operation, orange square is a consistency operation, green x is an equivalence query, and the red star represents the the upper bound termination condition. We observe it is possible to terminate early, prior to reaching the upper bound. Each plot contains 100 individual paths through phase space.
Figure 12: Phase diagram termination plots for CraftWorld tasks 5 through 10. Teacher tests 500 samples per equivalence query. Left to right (top row) Task 5, 6, 7; left to right (bottom row) Task 8, 9, 10. The X-axis is number of known representative values in the hypothesis, the Y-axis is the number of states in the hypothesis. Legend: blue circle is a closure operation, orange square is a consistency operation, green x is an equivalence query, and the red star represents the the upper bound termination condition. We observe it is possible to terminate early, prior to reaching the upper bound. Each plot contains 100 individual paths through phase space.
0.A.5 Remap Algorithms
@@
0.A.6 Technical Proofs
We first establish the correctness of constructing a hypothesis deterministic finite automaton from a symbolic observation table via Algorithm 7, in Theorem 1 via Propositions 1 through 5. The proof strategies for Propositions 3 through 5 generally follow those of Angluin [5], with appropriate adjustments to account for a set of possible hypotheses, compared to just a single hypothesis.
Definition 11
Let $h_{k}=\langle Q_{k},q_{0,k},\Sigma^{I},\Sigma^{O},\delta_{k},L_{k}\rangle$ be a Moore machine for some integer $k$. Two Moore machines $h_{1}$ and $h_{2}$ are structurally identical if a bijection $B:Q_{1}\rightarrow Q_{2}$ exists, and the transition functions $\delta_{1}$ and $\delta_{2}$ are consistent with $B$. That is, $B(\delta_{1}(q_{0,1},s))=\delta_{2}(q_{0,2},s)$ for all $s\in(\Sigma^{I})^{*}$ where $\delta_{1}$ and $\delta_{2}$ are the extended transition functions.
Proposition 1
@@ 0.A.6 Technical Proofs
MakeHypothesis$(\langle S,E,T;\mathcal{C},\Gamma\rangle)$. All hypotheses in $\mathcal{H}$ are structurally identical to each other, and differ only by their labeling function.
Proof
We first show that all hypotheses in $\mathcal{H}$ must be structurally identical by construction, but have different $\hat{L}$ functions. Identical structure for all hypotheses in $\mathcal{H}$ can be shown by observing that lines 1-3 of Algorithm 7 are the same for each possible hypothesis in $\mathcal{H}$ for a given $\langle S,E,T;\mathcal{C},\Gamma\rangle$ input. Line 1 is the construction for the set of states. This procedure is identical for all hypotheses in $\mathcal{H}$, so all hypotheses in $\mathcal{H}$ have the same set of states. Bijection between sets of states for any pair of hypotheses is satisfied by the identity bijection function. Furthermore, the transition function construction specified on Line 3 indicates that all transitions functions for all hypotheses within $\mathcal{H}$ are identical. Since the transition functions are identical, and since bijection between states is satisfied by the identify function, we have that $B(\delta_{k}(q_{0,k},s))=\delta_{k}(q_{0,k},s)$ by identity bijection, and furthermore, $\delta_{k}(q_{0,k},s)=\delta_{k^{\prime}}(q_{0,k^{\prime}},s)$ by Line 3 construction, since $\delta_{k}$ and $\delta_{k^{\prime}}$ are identical. This shows that all hypotheses in $\mathcal{H}$ are structurally identical.
Next, we show that each hypothesis in $\mathcal{H}$ has a unique labeling function. Note that in constructing the hypotheses, the only differences in the output hypothesis are due to lines 4 and 5. Lines 4 and 5 together select a satisfying solution for the free variables in $\langle S,E,T;\mathcal{C},\Gamma\rangle$, subject to the constraints $\mathcal{C}$. Therefore, each unique hypothesis in $\mathcal{H}$ is corresponds to a unique satisfying solution; the mapping between unique satisfying solutions and unique hypotheses is bijective.∎
Proposition 2
States in a Moore machine are represented by sets of sequences. Specifically, a mapping can always be constructed from sets of sequences to states, given an initial state.
Proof
We proceed by showing that a bijection always exists, simply by proposing a valid bijection for every case in an inductive argument. Suppose we have a finite alphabet $\Sigma$, and a Moore machine, with set of states $Q$, initial state $q_{0}\in Q$, and transition function $\delta:Q\times\Sigma\rightarrow Q$. We define the initial state $q_{0}$ of the Moore machine to correspond with the empty sequence $\varepsilon$ which has length 00, so in fact we have $q_{0}=\delta(q_{0},\varepsilon)$, and for any $q\in Q$, $q=\delta(q,\varepsilon)$. The transition function can be extended recursively to $\delta:Q\times\left(\Sigma^{*}\right)\rightarrow Q$ by observing that $\delta(q_{0},\sigma\cdot\omega)=\delta(\delta(q_{0},\sigma),\omega)$, where $\sigma$ is a sequence of length of 1, and $\omega$ is a sequence of length of at least $1$. Similarly, $\delta(q_{0},\omega\cdot\sigma)=\delta(\delta(q_{0},\omega),\sigma)$. Based on this extended transition function, we can now consider the set $Q^{\prime}=\{\delta(q_{0},\sigma)|\sigma\in\Sigma\}$ to be a subset of $Q$.
@@
Proof
We utilize the inductive proof of Lemma 2 from Angluin [5]. For the case of $s=\varepsilon$, with length 00, we have $\delta(q_{0},\varepsilon)=q_{0}=\textbf{row}(\varepsilon)$, which is true by definition. Now, let us assume that $\delta(q_{0},s)=\textbf{row}(s)$ holds for all $s\in(S\cup(S\cdot\Sigma))$ with lengths no greater than $k$. Let $s^{\prime}\in(S\cup(S\cdot\Sigma))$ be a sequence of length $k+1$ such that $s^{\prime}=s\cdot\sigma$. If $s^{\prime}\in S$, then $s\in S$ because $S$ is prefix closed. If $s^{\prime}\in S\cdot\Sigma$, then $s\in S$. We can then show that $\delta(q_{0},s^{\prime})=\delta(q_{0},s\cdot\sigma)=\delta(\delta(q_{0},s),\sigma)=\delta(\textbf{row}(s),\sigma)=\textbf{row}(s\cdot\sigma)=\textbf{row}(s^{\prime})$. This sequence of equalities follows as specified in Lemma 2 from Angluin [5].∎
Proposition 4
The entries of a unified, closed, and consistent symbolic observation table correspond to sequence classification consistent with constraints $\mathcal{C}$. Specifically, for a unified, closed, and consistent $\langle S,E,T;\mathcal{C},\Gamma\rangle$, let $\Sigma^{*}$ be partitioned into at most $k\leq|\Sigma^{O}|$ disjoint sets $F_{1},F_{2},...,F_{k}$, and let $\Sigma^{O}_{k}$ be a specific subset of $k$ distinct elements of $\Sigma^{O}$. A sequence $s\in\Sigma^{*}$ is a member of $F_{j}$ if and only if it is classified as $\sigma_{j}\in\Sigma^{O}_{k}$. The hypothesis output by MakeHypothesis$(\langle S,E,T;\mathcal{C},\Gamma\rangle)$ satisfies $\delta(q_{0},s\cdot e)\in F_{j}$ if and only if $T(s\cdot e)=\sigma_{j}$ for every $s\in(S\cup(S\cdot\Sigma))\text{ and every }e\in E$.
Proof
We adapt the inductive proof of Lemma 3 from Angluin [5], but make appropriate adjustments for sequence classification. Let $s\in(S\cup(S\cdot\Sigma))$ and $e=\varepsilon$ be the base case. Clearly, $\delta(q_{0},s\cdot\varepsilon)=\delta(q_{0},s)=\textbf{row}(s)$ as shown previously. If $s\in S$, then $\textbf{row}(s)\in F_{j}$ if and only if $T(s)=\sigma_{j}$. If $s\in S\cdot\Sigma$, then $\textbf{row}(s)\in\hat{Q}$, since the symbolic observation table is closed. A $\textbf{row}(s^{\prime})\in\hat{Q}$ is in $F_{j}$ if and only if $T(s^{\prime})=\sigma_{j}$.